mfkiddddddddd:

the ideal of a particular metal rod is 20.5cm. the measured length may vary from the ideal length by at most 0.045cm. what is the range of acceptable lengths for the rod A. x_<20.455 or x_>20.545 B. 20.455_> x >_20.545 C. 20.455<_ x <_20.545 D. 20.455 _> x <_20.545

LilChris:

The minimum range = 20.5 - 0.045 = 20.455 The maximum range = 20.5 + 0.045 = 20.545 The range is between 20.455 and 20.545 Writing this as inequality 20.455 ≤ x ≤ 20.545 brainly.com

mfkiddddddddd:

you are the best bro.